### v1

# RegNLP in Action: Facilitating Compliance Through Automated Information Retrieval and Answer Generation

Abstract
Regulatory documents, issued by governmental regulatory bodies, establish rules, guidelines, and standards that organizations must adhere to for legal compliance. These documents, characterized by their length, complexity and frequent updates, are challenging to interpret, requiring significant allocation of time and expertise on the part of organizations to ensure ongoing compliance.Regulatory Natural Language Processing (RegNLP) is a multidisciplinary subfield aimed at simplifying access to and interpretation of regulatory rules and obligations. We define an Automated Question-Passage Generation task for RegNLP, create the ObliQA dataset containing 27,869 questions derived from the Abu Dhabi Global Markets (ADGM) financial regulation document collection, design a baseline Regulatory Information Retrieval and Answer Generation system, and evaluate it with RePASs, a novel evaluation metric that tests whether generated answers accurately capture all relevant obligations and avoid contradictions.
RegNLP in Action: Facilitating Compliance Through Automated Information Retrieval and Answer Generation
Tuba Gokhan1, Kexin Wang2, Iryna Gurevych1,2, Ted Briscoe1 1Mohamed Bin Zayed University of Artificial Intelligence (MBZUAI) 2Ubiquitous Knowledge Processing Lab (UKP Lab) Department of Computer Science and Hessian Center for AI (hessian.AI) Technical University of Darmstadt
1 Introduction
Regulatory documents are formal quasi-legal texts issued by governmental regulatory bodies that set out rules, guidelines, and standards that organizations must follow to ensure compliance. These documents typically cover a wide range of topics, from environmental and financial compliance to workplace safety and data protection. Characterized by their complex and precise language, they are designed to cover all potential scenarios, but can be challenging to interpret without specialized knowledge. Additionally, regulatory documents are frequently updated to reflect new laws, technological changes, or shifts in societal norms, necessitating continuous monitoring and adaptation by organizations to meet new compliance requirements. Compliance with these regulations requires considerable investment in terms of time and expertise. For example, between 1980 and 2020, the U.S. public and private sectors spent an estimated 292.1 billion hours complying with 36,702 regulations, accounting for approximately 3.2% of total annual working hours Kalmenovitz (2023). Moreover, compliance errors can lead to severe penalties, as, for example, demonstrated by the recent C\$7.5 million fine levied against the Royal Bank of Canada by FINTRAC for failing to report suspicious transactionsBloomberg (2023). Regulatory Natural Language Processing (RegNLP) is a recent multidisciplinary subfield designed to simplify access to and interpretation of regulatory rules and obligations, reducing errors and enhancing operational efficiency.
This paper makes several significant contributions to RegNLP:
  1. 1.
    Automated Question-Passages Generation: We introduce a framework that employs Large Language Model (LLM) agents to automatically generate questions, with Natural Language Inference (NLI) integrated at the validation stage. This framework can be applied to specialized regulatory document collections to create datasets for the development and evaluation of question-answering and passage retrieval systems.
  2. 2.
    ObliQA: We have developed a multi-document, multi-passage QA dataset  [^1] https://github.com/RegNLP/ObliQADataset specifically developed to facilitate research in RegNLP. It consists of 27,869 questions and their associated source passages, all derived from the full collection of regulatory documentation provided by Abu Dhabi Global Markets (ADGM) [^2] https://www.adgm.com/, the authority overseeing financial services in the UAE’s free economic zones.
  3. 3.
    Regulatory Information Retrieval and Answer Generation: We introduce a task that first involves passage retrieval given a specific regulatory query, identifying all relevant obligations distributed across all regulatory documents. Following this, the task requires generating comprehensive answers based on the retrieved passages. These answers must synthesize information from multiple sources, ensuring that all pertinent obligations are included to fully address the regulatory query. The goal is to create responses that are both complete and concise, providing clear guidance without overwhelming the user with extraneous information.
  4. 4.
    Regulatory Passage Answer Stability Score, We develop RePASs a novel reference-free metric aimed at evaluating the quality of generated answers in regulatory compliance contexts. RePASs is tailored to the complex nature of regulatory texts, ensuring that all relevant obligations are accurately captured, fully reflected in the answers, ideally devoid of contradictions.
This paper is structured as follows. Section 2 discusses existing work on RegNLP. Section 3 describes the creation of the ObliQA dataset, and Section 4 introduces the RePASs evaluation metric. Section 5 describes our 2-step baseline Regulatory Information Retrieval and Answer Generation (RIRAG) system, which first retrieves relevant passages for a given query and secondly generates an answer from these passages, and presents our evaluation of both steps. Section 6 concludes the paper and provides directions for future work. Section Limitations addresses the limitations of this work.
2 Related Work
2.1 Research on Regulatory NLP
Recent RegNLP research has advanced significantly using various methodologies and datasets. Lau et al. (2005) utilized information retrieval and extraction on accessibility regulations from the US and Europe, formatted into unified XML for semi-structured data handling. Kiyavitskaya et al. (2008) employed the Cerno framework for semantic annotation to automate extracting rights and obligations from U.S. HIPAA and Italian laws. Chalkidis et al. (2018) introduced a hierarchical BiLSTM model for extracting obligations from legal contracts, demonstrating superior performance over traditional models. Nair et al. (2018) developed a pipeline for annotating global import-export regulations, enhancing compliance workflows. Chalkidis et al. (2021) proposed a two-step document retrieval process for EU/UK legislative compliance using BERT-based models, handling long document queries. Abualhaija et al. (2022) presented an automated question-answering system using BERT, achieving high accuracy in identifying relevant texts from European regulations, including the General Data Protection Regulation.
These studies highlight the potential of recent NLP techniques and tools for improving the retrieval and extraction of regulatory information. Our work extends RegNLP by firstly developing a novel dataset in the financial domain from the complete document collection of a single regulatory authority, and secondly defining a novel regulatory question-answering (QA) task and evaluation metric.
2.2 Synthetic Data Generation with LLMs
Recent research studies have demonstrated the versatile capabilities of LLMs in enriching QA datasets across diverse domains. Alberti et al. (2019) pioneered methods for generating synthetic QA corpora, using roundtrip consistency for validation and achieving state-of-the-art performance on tasks such as SQuAD2 and NQ. Maatouk et al. (2023) introduced a zero-shot learning approach for neural passage retrieval, using synthetic question generation and hybrid term-neural models to enhance retrieval performance without extensive domain-specific data. Abdullin et al. (2023) focused on synthetic dialogue dataset generation, automating dialogues for linear programming problems using LLMs, thereby aiding in training conversational agents. Maatouk et al. (2023) developed TeleQnA, specifically evaluating LLMs’ telecommunications knowledge, while Tihanyi et al. (2024) introduced CyberMetric for assessing cybersecurity knowledge. These datasets enable evaluation and further training of LLMs to enhance their domain-specific capabilities. Additionally, multilingual question generation for health education, as explored by Ackerman and Balyan (2024), highlights the effectiveness of LLMs in diverse linguistic contexts. Collectively, these studies underscore the transformative potential of LLMs in refining QA datasets tailored to specific domains, fostering the development of more specialized and effective AI tools.
2.3 Retrieval-Augmented Generation
RegNLP can benefit from advances in retrieval-augmented generation (RAG) Lewis et al. (2020) and other related technologies to simplify access to regulatory documents and improve compliance. Regulatory documents, known for their complexity and frequent updates, pose unique challenges that can be mitigated through methods that leverage the strengths of NLP and information retrieval systems. RAG has significantly improved the accuracy, efficiency, and trustworthiness of LLMs by integrating external, contextually relevant and up-to-date information. Notable approaches include: Self-RAG Asai et al. (2023) enhances response quality by incorporating self-reflection mechanisms; RAFT Zhang et al. (2024) optimizes domain-specific knowledge retrieval by filtering out irrelevant data and improving reasoning chains; RobustRAG Xiang et al. (2024) and RQ-RAG Chan et al. (2024) tackle challenges such as retrieval corruption and ambiguous queries; PipeRAG Jiang et al. (2024a) improves performance by reducing retrieval latency. However, these studies do not consider regulatory documents so we are interested in testing the ability of RAG methods for solving the QA task for regulatory questions.
3 ObliQA: Obligation Based Question Answering Dataset
The ObliQA - The Obligation Based Question Answering Dataset creation process involves three main steps: Data Collection, Question Generation, and Question-Passages Validation via NLI, as illustrated in Figure 1.
Figure 1: AI-driven Framework for Generating Question-Passage Pairs to Create ObliQA: Obligation-based Question-Answering Dataset for Regulatory Compliance. This flowchart presents the process from data collection and standardization to question generation and NLI-based validation for constructing a synthetic dataset.
#Question 1 Passage 2 Passages 3 Passages 4 Passages 5 Passages 6 Passages
Group 1 7988 7988
Group 2 7254 5432 1822
Group 3 5365 3385 1434 546
Group 4 4558 3057 1102 305 94
Group 10 2704 1325 678 345 174 121 61
Total 27869 21187 5036 1196 268 121 61
Table 1: Distribution of Questions by Number of Passages in the ObliQA Dataset. Groups indicate the number of input passages.
#Question 1 Passage 2 Passages 3 Passages 4 Passages 5 Passages 6 Passages
train 22295 16946 4016 975 202 100 56
test 2786 2126 506 105 36 9 4
development 2888 2215 514 116 30 12 1
Table 2: Distribution of Questions in the ObliQA Dataset Across Training, Testing, and Development sets , categorized by the number of associated passages.
3.1 Data Collection
The regulatory documents provided by ADGM are in .docx format, each ranging from 30 to 100 pages and written in a broadly legalistic style. These documents feature complex internal structures, including numerous (sub)subsections, numbered clauses, and cross-references within and across documents. In total, there are 40 documents comprising approximately 640,000 words.
[…]
Structuring the Data: Subsequently, the standardized .txt files are converted into a structured JSON format for data manipulation and analysis. Below is an example of the JSON structure:
"ID": "2230472f-a9d4-4b81-842f-964c0653f8e7",
"DocumentID": 20,
"PassageID": "1.1",
"Passage": "This Guidance has been produced to help legal practitioners in ADGM to understand the intention behind the Application of English Law Regulations 2015 (the ’Application Regulations’) and how English law has been implemented in ADGM..."
The PassageID is used to represent the hierarchical information of the passages within the documents.
3.1.1 Question Generation
Regulatory documents contain extensive descriptions and titles, which are filtered out in the question generation process to focus solely on essential content.
Input Definition:
Inputs for Single Passage Questions: Inputs for single passage questions isolate one specific source passage per question, simplifying the context and enhancing the relevance of the generated question.
Inputs for Multi-Passage Questions: For broader queries, inputs encompass multiple source passages. Sixty-nine distinct topics related to finance and regulatory compliance (see subsection A.1 for a sample) guide the aggregation of related regulations into meaningful clusters. These clusters then form the basis for multi-passage questions by randomly grouping a specified number of rules per question.
GPT-4 Generator: The gpt-4-turbo-1106 model is employed to generate both questions and answers.
For single passage questions, the prompt is:
"Your task is to generate realistic and applied questions that pertain to the provided regulatory or compliance material. Ensure that the context implicitly contains the answer to the question."
For multi-passage questions, the complexity increases, and the prompt is adjusted accordingly:
"Imagine you are preparing a list of questions that a compliance officer at a company would ask regulatory authorities like the ADGM to clarify their understanding of compliance requirements and to ensure adherence to regulatory standards. Each question should be direct, practical, and structured to obtain detailed information on specific regulatory rules. Avoid formulations that imply the company is describing its own processes, and instead frame questions that seek clarity on regulatory expectations."
3.2 Question-Passages Validation via NLI
Validating the generated question-passage pairs using Natural Language Inference (NLI) involves assessing the semantic relationship between each question and its corresponding source passages to ensure alignment. The validation process uses the nli-deberta-v3-xsmall model with the passage set as the premise and the question set as the hypothesis. For each passage associated with a question, NLI is applied to evaluate the relationship. The validation results are categorized as follows.
Entailment: The passages entail the questions being retained.
Contradictions: Passages that contradict the questions are eliminated.
Neutral: Passages with neutral scores are retained if they are closer to the entailment score but eliminated if they are closer to the contradiction score.
The validation process ensures that only passages with strong semantic alignment to the questions are included. This step results in a dataset where questions are linked to passages that provide relevant and accurate information.
The final dataset consists of approximately 21k questions linked to single passages and 5k questions linked to multiple passages. The distribution of questions among different groups, indicating the number of input passages, is detailed Table 1. The further distribution between the training, testing and development sets is shown in Table 2. The structure of the regulatory documents and the dataset can be found in the ObliQA Dataset Repository  [^3] https://github.com/RegNLP/ObliQADataset. A sample of the dataset is provided in Appendix A.2.
4 RePASs: Regulatory Passage Answer Stability Score
We propose a reference-free evaluation metric, RePASs, to evaluate generated answers in regulatory compliance contexts. Inspired by recent work on answer attribution to specific sources Bohnet et al. (2023); Yue et al. (2023), this metric evaluates answers based on three key aspects:
  • Every answer sentence must be supported by a sentence in the source passage(s).
  • The answer must not contain any sentences that contradict the information in the source passage(s).
  • The answer must cover all the obligations present in the source passages, meaning that all critical regulatory obligations should be reflected in the answer.
This metric prefers answers that are not only accurate but also comprehensive in covering all relevant regulatory obligations. The metric includes the following steps:
Step 1: Generation of Entailment and Contradiction NLI Pair Matrices
We generate an NLI pair matrix by comparing each passage sentence (premise) with each answer sentence (hypothesis) using the cross-encoder/nli-deberta-v3-xsmall He et al. (2021a). The model provides three outputs for each pair: probabilities for entailment, contradiction, and neutrality. These probabilities are organized into two matrices: the entailment matrix, which contains the entailment probabilities for each passage-answer sentence pair, and the contradiction matrix, which stores the contradiction probabilities for each sentence pair.
Step 2: Calculation of Entailment and Contradiction Scores
We reduce the size of the entailment and contradiction matrices by selecting the highest probability for each answer sentence across the passage sentences. Matrix reduction ensures that only the most relevant support from or conflict with the source passage is considered. The entailment score is computed by averaging the highest entailment probabilities across all answer sentences. Similarly, the contradiction score is obtained by averaging the highest contradiction probabilities.
$E_{s}=\frac{1}{N}\sum_{i=1}^{N}\max_{j}P_{\text{entailment}}(p_{j},a_{i})$ (1)
$C_{s}=\frac{1}{N}\sum_{i=1}^{N}\max_{j}P_{\text{contradiction}}(p_{j},a_{i})$ (2)
Where $E_{s}$ is the Entailment Score, and $C_{s}$ is the Contradiction Score. $N$ represents the total number of sentences in the generated answer. $P_{\text{entailment}}(p_{j},a_{i})$ is the probability that the $i$-th sentence in the answer ($a_{i}$) is entailed by the $j$-th sentence in the source passage ($p_{j}$). The $\max$ function selects the highest entailment probability for each answer sentence ($a_{i}$) across all sentences in the source passage ($p_{j}$). Similarly, $\max_{j}$ selects the highest probability for each answer sentence across all sentences in the source passage.
Step 3: Calculation of Obligation Coverage Score
The obligation coverage score assesses how accurately the generated answer reflects the obligations present in the source passage(s). First, we create a dataset using an automated pipeline powered by the GPT-4-turbo-1106 model. This pipeline extracted both obligation and non-obligation sentences from regulatory documents. The sentences are verified and labeled as obligations or non-obligations. The dataset was subsequently used to fine-tune a classifier model based on LegalBERT. Chalkidis et al. (2020) to define the obligatory sentences. Full details on the dataset creation and fine-tuning methods are available at https://github.com/RegNLP/ObligationClassifier. Both the source passages and the generated answers are tokenized into individual sentences, enabling sentence-level evaluation.
For each obligation sentence detected in the source passage, the system verifies its coverage in the generated answer. This is done by comparing the obligation sentences with those in the answer using the NLI model microsoft/deberta-large-mnli He et al. (2021b). If any sentence in the answer demonstrates an entailment score exceeding 0.7 when compared to an obligation sentence from the source passage(s), the obligation is considered covered. The obligation coverage score is calculated as the ratio of the number of covered obligations to the total number of obligations present in the source passage(s).
The formula for the Obligation Coverage Score is given by:
$OC_{s}=\frac{1}{M}\sum_{k=1}^{M}\mathbb{1}\left(\max_{l}P_{\text{entailment}}(o_{k},a_{l})>0.7\right)$ (3)
Where $OC_{s}$ represents the Obligation Coverage Score, and $M$ is the total number of obligation sentences in the source passage. $P_{\text{entailment}}(o_{k},a_{l})$ denotes the probability that the $k$-th obligation sentence in the source passage ($o_{k}$) is entailed by the $l$-th sentence in the answer ($a_{l}$). The $\max_{l}$ function selects the highest entailment probability for each obligation sentence across all sentences in the answer. The indicator function $\mathbb{1}$ returns 1 if the highest entailment score exceeds 0.7, meaning the obligation is considered covered.
Step 4: Calculation of the Regulatory Passage Answer Stability Score
The final step integrates the three components into a single composite score. The score penalizes contradictions and normalizes the result between 0 and 1.
$RePASs=\frac{\text{E}_{s}-\text{C}_{s}+\text{OC}_{s}+1}{3}$ (4)
To evaluate RePASs, we tested it on a set of genuine questions using question-related obligations and answers provided by ADGM experts. Table 3 presents performance of RePASs on these gold standard questions, relevant passages and expert answers. The high scores obtained suggest that RePASs aligns well with expert judgements and the result provides an indicative upper bound for the performance of an automated system.
Es Cs OCs RePASs
Real-Case 0.837 0.260 1.000 0.859
Table 3: Performance of RePASs on the gold standard data from ADGM.
5 Regulatory Information Retrieval and Answer Generation Task
The Regulatory Information Retrieval and Answer Generation (RIRAG) task we have defined breaks the task down into two distinct steps which are evaluated independently. It begins with a passage retrieval step that locates all obligations relevant to a given question within the regulatory documents. This is followed by the answer generation step, which aims to synthesize the retrieved data into precise and informative responses that comprehensively address the regulatory query. In this section, we describe a baseline approach for RIRAG, setting a standard that will support future advances and improvements in the field.
5.1 Passage Retrieval
We experiment with six retrieval models, which are (1) BM25 Robertson et al. (1994): the traditional lexical-based model; (2) DRAGON+ Lin et al. (2023): the State-of-the-Art (SotA) single-vector dense retriever model fine-tuned on MS MARCO Nguyen et al. (2016); (3) SPLADEv2 Formal et al. (2021): the SotA neural sparse retriever fine-tuned on MS MARCO; (4) ColBERTv2: the SotA multi-vector dense retriever model finetuned on MS MARCO; (5) NV-Embed-v2 Lee et al. (2024): the SotA single-vector dense retriever model fine-tuned on multiple text-embedding datasets, achieving the best [^4] As of Aug 30, 2024. performance on Massive Text Embedding Benchmark (MTEB) Muennighoff et al. (2023); (6) BGE-EN-ICL Xiao et al. (2023): the SotA single-vector dense retrieval model fine-tuned on multiple text-embedding datasets, achieving the second-best [^5] As of Aug 30, 2024. performance on MTEB.
DRAGON+, SPLADEv2 and ColBERTv2 represents the three main architectures for neural retrieval, achieving SotA effectiveness with limited training resources and parameters (all around 110M parameters). NV-Embed-v2 and BGE-EN-ICL are the best-performing dense retrievers when the training resources are extensive and the parameters are largely scaled up (7.1B and 7.9B parameters, respectively). Additionally, BGE-EN-ICL supports In-Context Learning (ICL) Brown et al. (2020) for text embeddings, providing potentials in obtaining task-adapted model with a low cost via adding example demonstration in the model input.
Since the retrieval task in RIRAG asks to retrieve passages from a corpus of long documents, we also consider it as a Document-Aware Passage Retrieval (DAPR) Wang et al. (2024) task to jointly model the document context and passage retrieval. To this end, we apply rank fusion to linearly fuse the passage ranking by the neural or BM25 retrievers and the document ranking by the BM25 retriever. The details are available in Appendix A.4. The counterpart without rank fusion is denoted as passage-only.
For evaluating the retrieval module in RIRAG, we use recall@10 as the main evaluation metric, as we rely on the retrieval module to cover the relevant information as much as possible while leaving the burden of noise filtering to the answer-generation module. This approach is in line with LongRAG Jiang et al. (2024b). To additionally assess the ranking ability within the top-10 passages, MAP@10 (Mean Averaged Precision) is also used for diagnosis purposes.
For the experimental setup, we truncate the queries and the passages to 512 tokens. In rank fusion, we fuse the ranking of top-100 passages and the ranking of all the documents. The fusion weight on the document side is set to 0.1. The ranking scores are normalized (cf. Equation 5) with respect to the top-100 passages into the range of $[0,1]$ for filtering the passages conveniently in the answer-generation step. For BGE-EN-ICL, we consider the scenarios of zero-shot, 1-shot, 3-shot and 5-shot. The ICL examples are obtained via random sampling from the training set of ObliQA. The instructions/query prefix [^6] These models do not support instructions for the passage input. in NV-Embed-v2 and BGE-EN-ICLL are “Given a question, retrieve passages that answer the question” and “Given a financial-regulation query, retrieve relevant passages that answer the query”, respectively.
5.2 Answer Generation
The answer generation process begins once 10 relevant passages have been retrieved for each query from the passage retrieval task. In our post-retrieval stage, we implement a score-Based filtering approach with a threshold of 0.2 to identify significant drops in relevance between consecutive passages. We also enforce a minimum score criterion of 0.7, ensuring that only the most relevant passages are considered for generating answers. Using these filtered passages, we use the gpt-4-turbo-1106 model to generate comprehensive answers. The model follows a tailored prompt designed to simulate the role of a regulatory compliance assistant, integrating all relevant obligations and best practices from the passages into a cohesive response. The prompt reads:
"You are a regulatory compliance assistant. Provide a detailed answer for the question that fully integrates all the obligations and best practices from the given passages. Ensure your response is cohesive and directly addresses the question. Synthesize the information from all passages into a single, unified answer."
For evaluation, the responses are organized in JSON format, as shown in the Appendix A.3.
5.3 Results
5.3.1 Evaluation Dataset
Our method for the regulatory information retrieval and answer generation task is evaluated using a dataset of 446 question-passage pairs, which have been validated by ADGM experts. The detailed breakdown of the evaluation dataset is illustrated in Table 4.
# Passages per Question # Questions
1 Passage 182
2 Passages 174
3 Passages 69
4 Passages 13
5 Passages 4
6 Passages 4
Table 4: Distribution of questions by number of passages in the evaluation dataset.
5.3.2 Passage Retrieval Results
The results of the retrieval task are shown in Table 5. We find the models of DRAGON+, SPLADE and ColBERTv2, which are fine-tuned with limited training data (i.e. MS MARCO only) and limited parameters generalize poorly, achieving worse recall@10 and MAP@10 than simple BM25 most of the time. ColBERTv2, the best model in this group, outperforms BM25 by 1.8 points MAP@10 and 0.3 points recall@10. On the other hand, the models of NV-Embed-v2 and BGE-EN-ICL, which are fine-tuned with more training data and more parameters, outperform BM25 significantly in all the settings. Compared with BM25, recall@10 and MAP@10 are improved by up to 4.9 points and up to 2.9 points, respectively. These results show that scaling the retrievers with more training data and parameters is necessary for good generalization to the retrieval task in RIRAG. We find ICL is important for adapting the neural models to this new task, while the advantage of including more examples in the demonstration is very marginal. Lastly, we find rank fusion, which considers the document context during passage retrieval, improves the performance only marginally. We assume this is because the query generation process only considers the target passage as the input, resulting in fairly self-contained query-evidence pairs.
Model Passage-only Rank fusion
R@10 M@10 R@10 M@10
BM25 64.2 50.9 64.2 51.0
DRAGON+ 61.4 46.3 61.9 46.3
SPLADE 64.2 49.6 64.1 49.5
ColBERTv2 64.5 52.7 64.6 52.7
NV-Embed-v2 67.2 52.2 67.0 52.0
BGE-EN-ICL (0-shot) 66.5 51.9 66.6 52.0
BGE-EN-ICL (1-shot) 68.3 53.8 68.5 53.7
BGE-EN-ICL (3-shot) 68.3 53.8 68.7 54.0
BGE-EN-ICL (5-shot) 69.1 53.7 69.2 53.9
Table 5: Results of the retrieval task on the evaluation dataset. R@10 and M@10 represent Recall@10 and MAP@10, respectively.
5.3.3 Answer Generation Results
Method Es Cs OCs RePASs
BM25(passage-only)+GPT-4 0.769 0.238 0.220 0.583
BM25(rank fusion)+GPT-4 0.775 0.240 0.204 0.579
Table 6: Results of the answer generation task using RePASs on the evaluation dataset. Es, Cs, OCs and RePASs represent Entailment, Contradiction, Obligation Coverage and RePAS score, respectively.
The Table 6 illustrates the performance differences between the methodologies BM25(passage-only)+GPT-4 and BM25(rank fusion)+GPT-4 used in our answer generation task. Both methods show comparable performance in terms of Es and Cs. However, BM25(passage-only)+GPT-4 exhibits a slight advantage in the OCs, leading to a higher overall RePASs.
6 Conclusion
In conclusion, this paper advances the field of RegNLP by introducing innovative tools and methodologies aimed at improving the overall efficiency and precision of regulatory compliance. The automated question-passage generation framework and the ObliQA dataset significantly contribute to refining the retrieval and understanding of regulatory information. Furthermore, the regulatory information retrieval and answer generation task, supported by the novel RePASs evaluation metric, provides a robust framework for improving on our baseline approach.
In the future, the field of RegNLP could expand with the integration of new tasks. For instance, Summarization and Simplification methods could be used to develop tools to make complex regulatory texts simpler and more accessible to nonexperts. Automated Compliance Checking could leverage cross-document analysis techniques to improve regulatory adherence by comparing an organization’s internal compliance documents with regulatory documents. Regulatory Gap Analysis could focus on identifying vagueness, ambiguity or contradictions in regulations to aid in their refinement and ongoing compliance when regulations are added or updated. The continued development of these tasks will equip RegNLP to better serve the needs of regulatory bodies and regulated organizations alike, making compliance a more streamlined, reliable and efficient process.
Limitations
One significant limitation we encountered is the scarcity of publicly available datasets. Regulatory compliance often involves handling highly sensitive and confidential information, such as personal data or financial records, which presents challenges in terms of data privacy and security. To address this, we created a semi-synthetic dataset—ObliQA—using our Automated Question-Passage Generation Framework. During the validation phase, we faced difficulties stemming from multidisciplinary complexity and shortage of experts. However, experts evaluated a small subset of the dataset and confirmed that 99.7% of the generated questions met expectations, while 86.85% of the answer passages were deemed relevant to the questions.
Another key limitation is the lack of standardized formats for regulatory documents. Regulatory documents vary significantly across industries, regions, and even within the same organization, making it difficult to develop models capable of generalization. In some cases, even within the same regulatory authority, documents are prepared in different formats due to departmental variations. In this work, standardizing 40 documents required over 300 person-hours of effort. Additionally, the handling of tables and figures presented further challenges; tables were semi-automatically converted to text format, and figures were interpreted using the GPT-4 model.
Furthermore, the absence of comprehensive benchmarks is another limitation. There is no standardized benchmarking system available to evaluate the performance of RegNLP models across various regulatory domains, making it challenging to assess the effectiveness of different approaches or track progress in the field. Although we proposed the RePASs metric, we were unable to compare it with other models due to the lack of obligation coverage in existing benchmarks. Nevertheless, on genuine gold standard data from ADGM, RePASs demonstrated its suitability for RIRAG tasks.
Acknowledgement
We would like to express our gratitude to the Financial Services Regulatory Authority (FSRA)’s Applied Innovation & Research Team within ADGM for providing data and annotations. We also gratefully acknowledge ADGM’s support through a grant, enabling access to OpenAI GPT models via the Azure cloud, which contributed significantly to the technical development of this research.
Appendix A Appendix
A.1 Keyword-Topic List for ObliQA Dataset
topics_keywords = {
"AML Compliance": [
"money laundering", "compliance program", "regulatory requirements"
"Anti-Money Laundering": [
"AML","money laundering", "KYC", "know your customer", "financial crime", "terrorist financing",
"due diligence", "suspicious activity", "compliance program"
"Audit and Monitoring": [
"audit", "compliance monitoring", "internal audits", "external audits"
"Blockchain Technology": [
"blockchain", "blockchain technology", "smart contract", "tokenization"
"Blockchain-based Securities": [
"blockchain-based securities", "blockchain technology in securities"
"Virtual Asset Regulation": [
"virtual assets", "crypto assets", "digital asset regulation", "virtual asset service providers", "VASP", "crypto exchanges", "crypto custodians", "ICO regulations", "token classifications"
A.2 JSON Data Examples from the ObliQA Dataset
"QuestionID": "739921c1-385a-4735-a052-dee9fba73602",
"Question": "What are the key compliance indicators that a Fund Manager should monitor to confirm that a Passported Fund is being managed and operated within its constitutional framework and applicable ADGM legislation?",
"Passages": [
"DocumentID": 16,
"PassageID": "Part 3.6.(2)",
"Passage": "Each Reporting UAE Financial Institution shall establish and implement appropriate systems and internal procedures to enable its compliance with the Cabinet Resolution and these Regulations."
"DocumentID": 5,
"PassageID": "6.1.2",
[…]
"QuestionID": "8dc451a4-ec55-4fa5-abce-b0b8764a9338",
"Question": "Can the Regulator provide case studies or examples where the conduct of an Approved Person has been deemed compliant or non-compliant with the Principles, to help clarify the practical application of these rules?",
"Passages": [
"DocumentID": 6,
"PassageID": "PART 5.13.3.9.Guidance.3",
"Passage": "The onus will be on the Regulator to show that he is culpable, taking into account the standard of conduct required under the principle in question. In determining whether or not the particular conduct of an individual complies with the principles, the Regulator will take into account whether that conduct is consistent with the requirements and standards relevant to an individual’s role and the information available to him."
"DocumentID": 7,
"PassageID": "2.3.1.Guidance.3.",
[…]
A.3 JSON Data Example for Evaluation of Answer Generation Step
"QuestionID": "12345678-abcd-1234-efgh-ijklmnopqrst",
"Question": "What are the regulatory requirements for data encryption in financial services?",
"RetrievedPassages": [
"Financial institutions must encrypt sensitive data at rest and in transit. Key requirements include the use of AES-256 encryption for data at rest and TLS 1.2 or higher for data in transit."
"Answer": "Regulatory requirements for data encryption in financial services mandate using AES-256 encryption for data at rest and TLS 1.2 or higher for data in transit to ensure security of sensitive information.",
"RetrievedIDs": [
"9abc1234-def5-6789-ghij-klmnopqrstuv"
A.4 Rank Fusion
Rank fusion fuses the relevance scores from a BM25 retriever and a neural retriever. We compute the fusion as the convex combination of the normalized relevance scores Wang et al. (2021):
$s_{\mathrm{convex}}(q,p,d)$
$=\alpha\hat{s}_{\mathrm{BM25}}(q,p)+(1-\alpha)\hat{s}_{\mathrm{neural}}(q,d),$
where $\alpha\in[0,1]$ is the fusion weight, and $\hat{s}_{\mathrm{BM25}}$/$\hat{s}_{\mathrm{neural}}$ represents the normalized BM25/neural-retrieval relevance score, respectively. The normalization for a relevance score $s$ is calculated as:
$\hat{s}(q,c)=\frac{s(q,c)-m_{q}}{M_{q}-m_{q}},$ (5)
where $c$ represents the passage/document candidate, $m_{q}$ and $M_{q}$ are the min. and max. relevance scores of the top candidates for $q$, respectively. For any misaligned candidates, a zero score is taken for the candidate-missing side. In this work, the fusion is applied between the BM25 document rankings and the neural passage rankings of a certain cut-off.